# Adaptive Modality Balanced Online Knowledge Distillation for Brain-Eye-Computer based Dim Object Detection

### v2

@@ Abstract
Advanced cognition can be extracted from the human brain using brain-computer interfaces. Integrating these interfaces with computer vision techniques, which possess efficient feature extraction capabilities, can achieve more robust and accurate detection of dim targets in aerial images. However, existing target detection methods primarily concentrate on homogeneous data, lacking efficient and versatile processing capabilities for heterogeneous multimodal data. In this paper, we first build a brain-eye-computer based object detection system for aerial images under few-shot conditions. This system detects suspicious targets using region proposal networks, evokes the event-related potential (ERP) signal in electroencephalogram (EEG) through the eye-tracking-based slow serial visual presentation (ESSVP) paradigm, and constructs the EEG-image data pairs with eye movement data. Then, an adaptive modality balanced online knowledge distillation (AMBOKD) method is proposed to recognize dim objects with the EEG-image data. AMBOKD fuses EEG and image features using a multi-head attention module, establishing a new modality with comprehensive features. To enhance the performance and robust capability of the fusion modality, simultaneous training and mutual learning between modalities are enabled by end-to-end online knowledge distillation. During the learning process, an adaptive modality balancing module is proposed to ensure multimodal equilibrium by dynamically adjusting the weights of the importance and the training gradients across various modalities. The effectiveness and superiority of our method are demonstrated by comparing it with existing state-of-the-art methods. Additionally, experiments conducted on public datasets and system validations in real-world scenarios demonstrate the reliability and practicality of the proposed system and the designed method. The dataset and the source code can be found at: https://github.com/lizixing23/AMBOKD.
Index Terms:
object detection, EEG, multimodal learning, online knowledge distillation, adaptive modality balancing
@@ I Introduction
In recent years, considerable progress has been made in the field of computer vision, primarily due to rapid advancements in deep learning [1, 2]. Remarkable results in different visual tasks, such as object recognition, image generation, and visual localization, have been achieved by combining deep learning with high-performance computers, well-designed neural networks, and large datasets [3]. However, accurately and robustly detecting dim objects in aerial images remains a difficult task because of factors such as cluttered backgrounds, varying observing angles, and small object scales [4]. Furthermore, the performance of dim object detector is further restricted by the limited training samples of sensitive objects in aerial images.
Figure 1: Data extraction and processing procedure based on adaptive modality balanced online knowledge distillation (AMBOKD) method. See https://youtu.be/Yj01V8TBlUQ for an overview of the experiment and the AMBOKD method.
@@
A promising avenue to address the above limitations is offered by brain-computer interfaces (BCIs). This technique is able to decode brain activity in response to events, providing insights into human cognitive processes [5]. Leveraging the prior knowledge and advanced cognitive abilities of humankind, BCIs can compensate the shortcomings of comuputer vision systems and achieve more robust and accurate recognition capabilities. Existing studies have demonstrated the potential of using event-related potential (ERP) detection for target recognition tasks. [6, 7, 8]. By detecting and interpreting ERPs elicited from electroencephalogram (EEG) signals, these works have shown effectiveness in target recognition tasks across different scenarios [6]. However, due to the low noise of the EEG signals, the accuracy of EEG-based target recognition approaches still needs to be enhanced.
To overcome the inherent shortcomings of computer vision-based approaches and BCIs, a straightforward idea is to combine these two approaches. Recently, multimodal learning has emerged as a powerful tools to incorporate the advantages of computer vision and EEG modalities by ingratiating diverse data and essential information [9]. Researchers have investigated multimodal learning through diverse strategies using various structures such as early fusion [10] at the feature level and late fusion [11] employing logistic regression. However, most current studies mainly focus on feature fusion or loss function, without thoroughly considering the relationships between various modalities, thereby impeding the promotion of mutual learning and synergistic effects among modalities [12]. Thus, extracting critical information from multimodal data (i.e., visual and EEG data) and improving recognition performance remains an ongoing challenge.
Knowledge distillation (KD), a widely used technique for knowledge transfer, offers a solution to the aforementioned challenge. KD provides a paradigm where the student model can learn more information from the teacher model, rather than relying solely on the correctness of the true labels. This kind of method has similarities with human learning, deriving various algorithms in the field of multimodal learning, such as cross-modal [13] and multi-teacher [14] KD, which improves performance and generalization while reducing the model parameters [15]. In order to simplify the distillation procedure and circumvent the disadvantages associated with conventional KD in terms of training cost and pipeline complexity, online knowledge distillation (OKD) has been proposed to facilitate mutual learning between models in the end-to-end training process [16] [17]. Nevertheless, most current OKD approaches depend on a single modality and are restricted to supporting homogeneous networks. These approaches fail to address the imbalance issues arising from modal differences during the online learning process, lacking the ability to extract and integrate multimodal information.
In this study, a brain-eye-computer based object detection system is built for detecting dim object in aerial images under few-shot conditions, which is able to leverage the strengths of humans in rapid cognitive and computer vision in data processing. This system uses the region proposal networks to detect suspicious object region, evokes the ERP signal with the eye-tracking-based slow serial visual presentation (ESSVP) paradigm, and consturcts EEG-image data pairs with eye movement data and recognizes them. In particular, an adaptive modality balanced online knowledge distillation (AMBOKD) method is proposed to process the EEG-image data and recognize the target image. As shown in Fig. 1, the experimental procedure begins with the acquisition of EEG and image data, which are input into their respective encoders. The outputs of these encoders are fused by the fusion module, which is considered the third modality. Subsequently, the visual, EEG, and fusion modalities exchange roles between teacher and student models. For example, when visual modality acts as the student model, EEG modality and fusion modality act as the teacher models. This setup facilitates mutual learning and collective optimization using the OKD method. During mutual learning, the adaptive modality balancing (AMB) module is designed to facilitate parameter optimization for all modalities by dynamically balancing the influence weights and training optimization level of each modality. Consequently, the proposed approach is well suited for multimodal heterogeneous tasks, enabling the fusion of 2D EEG signals and 3D visual image data, leading to robust and efficient performance in dim object recognition.
@@
To the best of our knowledge, this study is the pioneering effort in addressing the challenges of dim object detection in aerial images using heterogeneous multimodal data. The main contributions of this study can be summarized as follows:
  • A brain-eye-computer based object detection system is established to obtain the subject’s attention region image and EEG data. This system is able to fully untilize the advantages of multi modalities, and detect the dim target in aerial images under few-shot conditions.
  • An AMBOKD method is proposed to fuse the multimodal data and enable simultaneous end-to-end mutual learning for target recognition. Under this method, an AMB module is introduced to adaptively balance the influence weights and training gradients of each modality, so as to ensure comprehensive parameter optimization.
  • A multimodal ESSVP dataset is built with 224×224 size RGB images and 59-channel EEG data. This dataset contains more than 13,000 paired samples of EEG and images, presenting promising prospects for the further development of multimodal learning methods in target detection tasks.
  • The superiority, robustness, and generalizability of our method is demonstrated by comparing it with state-of-the-art methods in a series of experiments. Our method not only improves the performance of the fusion model but also enhances the capabilities of both EEG and visual models, highlighting its potential in the field of multimodal learning.
@@
The reset of this paper is structured as follows: Section II discusses the related work about this study. Section III presents the design of the built system and the ESSVP dataset in detail. Section IV illustrates the proposed target recognition method. Section V presents the experimental results and analyzes the performance. Section VI draws conclusions and possible directions for future work.
@@ II-A Brain-Computer Based Object Recognition
Computer-based target recognition approaches has been fully developed because of the rapid development of deep learning. The well-known neural network algorithms commonly employed in this domain include ResNet [18], MobileNetV2 [19], VGG [20], DenseNet [21], and EfficientNet [22]. Notably, ResNet [18], which can learn residual mappings, enables the network to effectively capture and represent complex features, leading to high-accuracy classification performance. EfficientNet [22] focuses on achieving high performance with computational efficiency, achieved through a meticulously designed network architecture and scaling method. However, the existing algorithms still suffer from false alarms and missed detection problems. These algorithms are highly dependent on the data set due to the influence of factors such as the target environment, training data and noise.
@@
In recent years, EEG-based object recognition approaches have attracted considerable attention because of their ability to extract human cognition [23, 24]. These approaches depend on the analysis of ERP signals generated when subjects identify a target, providing enhanced robustness in intricate environments and minimizing the need for large amounts of image data [25, 7, 26]. The rapid serial visual presentation (RSVP) paradigm evokes ERP signals employed for target recognition by presenting visual image sequences as stimuli. For instance, Bigdely-Shamlo et al. [27] conducted an RSVP experiment where subjects were instructed to identify images containing airplanes at a frame rate of 12 Hz. They achieved high accuracies on 128-channel EEG data using independent component analysis (ICA), and reported promising results for single-trial classification. Lan et al. [7] proposed a multi-attention convolutional recurrent model for ERP detection, facilitating the identification of target images within RSVP sequences [28]. In addition, Fan et al. [26] introduced an asynchronous visual evoked paradigm (AVEP) based on RSVP, and proposed a deep learning method for detecting dim objects in satellite images. However, EEG signals are different on individual subjects and are prone to noise interference during signal extraction, thereby necessitating further improvements.
To harness the benefits of EEG and computer vision approaches, the fusion of EEG and image features has been investigated to accomplish efficient and robust object recognition. Barngrover et al. [29] developed a specific brain-computer Interface (BCI) system that addresses the challenge of target mine recognition inside-scan sonar images by fusing image and EEG signal features. Huang et al. [30] proposed a Bayesian HV-CV retrieval framework (BHCR) that combines human and computer vision using Bayesian approaches. The RSVP experimental paradigm was used to recognize targets, leading to effective retrieval of image databases. Minor et al. [31] introduced a multimodal neural network that combines EEG and image features. By fusing these modalities, they achieve high classification results in object recognition tasks. Previous studies have indicated the potential of integrating EEG and visual modalities to enhance the efficiency and robustness of recognition systems.
In line of this, our study aims to achieve dim object detection on aerial images by constructing the brain-eye-computer based object detection system with the ESSVP paradigm, which is build upon the RSVP and AVEP paradigms. This system detects the suspicious object region with the region proposal networks and employs eye-tracking technology to simultaneously extract the attention region and EEG signal fragment of the subject from the image. The proposed multimodal fusion approach outlined in this study then uses these two modal data sources as inputs.
@@ II-B Multimodal Learning
Modeling and analyzing data from various sensory modalities, including images, speech, text, and EEG signals, are key aspects of multimodal learning [32, 33, 34]. Significant attention has been garnered by multimodal learning in the fields of artificial intelligence and machine learning in recent years, resulting in remarkable progress [9, 35]. Currently, different multimodal applications, such as human-computer interaction [33, 36], natural language translation [37, 38], and computer vision [39, 34], are being applied in our lives.
Due to the development of deep learning, researchers propose novel loss functions or training approaches to optimize multimodal models. For instance, Nagrani et al. [40] introduced a transformer-based architecture that employs fusion bottlenecks for modality fusion at multiple layers. This method enables each modality to capture crucial information while efficiently sharing necessary information, leading to enhanced fusion performance and reduced computational cost. Du et al. [32] modeled the relationships between brain, visual, and linguistic features using multimodal deep generative models. They maximized both intra-modality and inter-modality mutual information regularization terms. Their approach addresses limitations such as the under-exploitation of multimodal knowledge and the scarcity of training data. Yang et al. [41] proposed a multimodal fusion approach for remote sensing image-audio retrieval tasks. They converted audio inputs to text and fused them with the text information to obtain a fusion representation. They optimized the common retrieval space using triplet loss, semantic loss, and consistency loss. Experimental findings on multiple datasets indicate the effectiveness of their approach. However, these methods primarily focus on feature fusion and loss function design when developing multimodal fusion algorithms, thereby overlooking the potential benefits of knowledge transfer and mutual learning between different modalities.
In this paper, we propose a dynamic hybrid fusion approach that integrates multimodal features and dynamically trains all branches in the model using an OKD method. This mutual learning approach for multimodal fusion aims to fully leverage the benefits of the visual and cognitive domains and extract crucial information effectively.
@@ II-C Knowledge Distillation
KD has attracted considerable attention because of its capacity to compress models. This technique involves training a small student model by imitating the output distribution of a large teacher model [15]. KD has been extensively employed in different fields, including computer vision [42, 43], natural language processing [44, 45], and emotion recognition [36]. These appoarches can be classified into offline distillation and online distillation according to the distillation schemes [46]. Offline distillation typically follws a two-stage training process where a trained teacher model guides the training of a student model. However, this approach is expensive and knowledge transfer is unidirectional. In contrast, online approaches enable end-to-end training, allowing for simultaneous learning of teacher and student models while reducing training [47].
@@
Current online knowledge distillation approaches have achieved significant progress with multimodal homogeneous data. For instance, Zhang et al. [47] introduced DML, demonstrating that student models can learn from each other through their predictions in deep mutual learning. Guo et al.[52] presented an OKD approach using collaborative learning with a weighted ensemble logit distribution. Li et al. [53] proposed an innovative embedded OKD approach that surpasses existing OKD approaches in image classification tasks. This approach leverages ensemble information, overall feature representations from peer networks, and logits to fully exploit the potential of networks. However, there remains a research gap in online knowledge distillation methods for multimodal heterogeneous data, due to the challenges in extracting and fusing heterogeneous data features.
To this end, we propose the AMBOKD method, which fully extracts features from both 2D EEG data and 3D image data, and treats the features output by the fusion model as a single modality for mutual learning. This method adaptively adjust the influences weights and the optimization level of each modality, leading to fully optimization of the parameters and surperior performance of the fusion model, thereby opening up new application scenarios for multimodal OKD.
@@ III System Design and Data Collection
In this section, we start by introducing the structure and data acquisition process of the brain-eye-computer based object detection system. Then, we verify the effectiveness of the employed RPN network. The designed ESSVP paradigm and its experimental environment are illustrated in detail. Lastly, we present the preprocessing process and the details of the ESSVP dataset.

### v3

@@ Abstract
Advanced cognition can be measured from the human brain using brain-computer interfaces. Integrating these interfaces with computer vision techniques, which possess efficient feature extraction capabilities, can achieve more robust and accurate detection of dim targets in aerial images. However, existing target detection methods primarily concentrate on homogeneous data, lacking efficient and versatile processing capabilities for heterogeneous multimodal data. In this paper, we first build a brain-eye-computer based object detection system for aerial images under few-shot conditions. This system detects suspicious targets using region proposal networks, evokes the event-related potential (ERP) signal in electroencephalogram (EEG) through the eye-tracking-based slow serial visual presentation (ESSVP) paradigm, and constructs the EEG-image data pairs with eye movement data. Then, an adaptive modality balanced online knowledge distillation (AMBOKD) method is proposed to recognize dim objects with the EEG-image data. AMBOKD fuses EEG and image features using a multi-head attention module, establishing a new modality with comprehensive features. To enhance the performance and robust capability of the fusion modality, simultaneous training and mutual learning between modalities are enabled by end-to-end online knowledge distillation. During the learning process, an adaptive modality balancing module is proposed to ensure multimodal equilibrium by dynamically adjusting the weights of the importance and the training gradients across various modalities. The effectiveness and superiority of our method are demonstrated by comparing it with existing state-of-the-art methods. Additionally, experiments conducted on public datasets and real-world scenarios demonstrate the reliability and practicality of the proposed system and the designed method. The dataset and the source code can be found at: https://github.com/lizixing23/AMBOKD.
ABBREVIATIONS
Abbreviation Full Form
AMBOKD Adaptive modality balanced online knowledge distillation
AMB Adaptive modality balancing
ANOVA Analysis of variance
BCIs Brain-computer interfaces
CE Cross-entropy
EEG Electroencephalogram
ERP Event-related potential
ESSVP Eye-tracking-based slow visual presentation
KD Knowledge distillation
OKD Online knowledge distillation
RPN Region proposal network
RSVP Rapid serial visual presentation
@@ I Introduction
In recent years, considerable progress has been made in the field of computer vision, primarily due to rapid advancements in deep learning [1]. Remarkable results in visual tasks, such as object recognition, image generation, and visual localization, have been achieved through deep learning, leveraging high-performance computing, well-designed neural networks, and large datasets [2]. However, accurately and robustly detecting dim objects in aerial images remains challenging due to cluttered backgrounds, varying observing angles, and small object scales [3]. Moreover, in target-unknown detection tasks, such as battlefield reconnaissance and disaster rescue, limited training samples from UAVs hinder accurate target detection by purely computer vision-based methods. Human judgment is often required to address these challenges.
@@
While manual methods are inefficient, brain-computer interfaces (BCIs) present a promising solution to these challenges. This technique is able to decode brain activity in response to events, leveraging the prior knowledge and advanced cognitive abilities of humankind. Existing studies have demonstrated the potential of using event-related potential (ERP) detection for target recognition tasks [4, 5]. By detecting and interpreting ERPs in electroencephalogram (EEG) signals, these works have shown effectiveness in target recognition tasks across different scenarios [6]. However, due to the high noise of the EEG signals, the accuracy of EEG-based target recognition approaches still needs to be enhanced.
Research demonstrates that computer vision systems effectively transform visual stimuli into discriminative semantic representations by leveraging large-scale imaging transformations and modeling intra-class variations [7]. In contrast, human visual processing achieves subconscious high-level feature extraction through the dynamic interplay between bottom-up stimulus-driven mechanisms and top-down attentional modulation [8]. Multimodal learning approaches integrate these feature representations, leveraging the complementary advantages of computer vision and EEG modalities [9]. However, current multimodal fusion approaches [10, 11, 12] typically rely on static weighting strategies, which assign fixed importance coefficients to different modalities throughout the training process. This fundamental assumption contradicts the empirical observation that modality reliability evolves dynamically, particularly evident in EEG-visual fusion where visual features typically converge faster than neural signals. Furthermore, existing knowledge distillation frameworks [13, 14, 15] often rely on unidirectional knowledge transfer patterns, neglecting the potential synergistic effects that could emerge from bidirectional cross-modal interactions. These inherent limitations become particularly detrimental in few-shot aerial detection scenarios, where the scarcity of training data amplifies the impact of suboptimal fusion strategies.
Figure 1: Brain-eye-computer based object detection system with our proposed AMBOKD method.
In this study, we develop a brain-eye-computer-based object detection system for dim objects in aerial images under few-shot conditions, leveraging human cognitive strengths and computer vision’s data processing capabilities. This system employs region proposal networks to detect suspicious object regions, elicits ERP signals using the eye-tracking-based slow serial visual presentation (ESSVP) paradigm, and constructs EEG-image data pairs with eye movement data for target recognition. As shown in Fig. 1, the proposed adaptive modality balanced online knowledge distillation (AMBOKD) method establishes a tri-modal framework where the fusion model operates as an independent modality, enabling mutual knowledge exchange with both visual and EEG streams. This framework dynamically adjusts modality-specific learning rates and cross-modal knowledge weights in real-time, ensuring balanced optimization across heterogeneous data representations.
@@
To the best of our knowledge, this study is the pioneering effort in addressing the challenges of dim object detection in aerial images using heterogeneous multimodal data, which holds great promise in fewshot scenarios such as military reconnaissance and disaster search and rescue. The main contributions of this study can be summarized as follows:
  • A novel brain-eye-computer based object detection system is established, by leveraging the complementary strengths of machine vision and human vision, achieving superior performance in detecting dim targets in aerial images.
  • An AMBOKD method is proposed to fuse heterogeneous multi-modal data, which significantly enhancing detection performance through online mutual learning with dynamic modality balancing.
  • The first publicly available ESSVP dataset is introduced, which incorporates temporally aligned EEG-visual pairs with aerial imaging characteristics, addressing a crucial gap in multimodal brain-computer vision research.
@@
The rest of this paper is structured as follows: Section II discusses the related work of this study. Section III presents the design of the built system and the ESSVP dataset in detail. Section IV illustrates the proposed target recognition method. Section V presents the experimental results and analyzes the performance. Section VI draws conclusions and possible directions for future work.
@@ II-A Brain-Computer Based Object Recognition
Computer-based target recognition approaches have been fully developed because of the rapid development of deep learning. The well-known neural network algorithms commonly employed in this domain include ResNet [16], MobileNetV2 [17], and EfficientNet [18]. However, the existing algorithms still suffer from false alarms and missed detection problems. These algorithms are highly dependent on the data set due to the influence of factors such as the target environment, training data and noise.
@@
In recent years, EEG-based object recognition approaches have attracted considerable attention because of their ability to extract human cognition [19]. These approaches depend on the analysis of ERP signals generated when subjects identify a target, providing enhanced robustness in intricate environments and minimizing the need for large amounts of image data [4, 20]. The rapid serial visual presentation (RSVP) paradigm evokes ERP signals employed for target recognition by presenting visual image sequences as stimuli. For instance, Lan et al. [4] proposed a multi-attention convolutional recurrent model for ERP detection, facilitating the identification of target images within RSVP sequences [21]. In addition, Fan et al. [20] introduced an asynchronous visual evoked paradigm (AVEP) based on RSVP, and proposed a deep learning method for detecting dim objects in satellite images. However, EEG signals are different on individual subjects and are prone to noise interference during signal extraction, thereby necessitating further improvements.
To harness the benefits of EEG and computer vision approaches, the fusion of EEG and image features has been investigated to accomplish efficient and robust object recognition. Barngrover et al. [22] developed a specific brain-computer Interface (BCI) system that addresses the challenge of target mine recognition inside-scan sonar images by fusing image and EEG signal features. Song et al. [23] employed a self-supervised framework to demonstrate the feasibility of learning image representations from EEG signals. Their extensive experiments affirm the biological plausibility of this approach, providing insights into human object recognition from temporal, spatial, spectral, and semantic perspectives. Quan et al. [7] utilized the rhesus monkey dataset [24] to analyze the similarities and differences between image representations derived from brain responses and those computed by deep convolutional neural networks, demonstrating the feasibility of EEG-visual fusion. Furthermore, they developed an EEG-visual fusion model to jointly learn brain responses and image representations from convolutional neural networks, validating the model’s performance on a self-collected dataset.
However, existing approaches mainly target salient objects or simple scenarios, neglecting to develop robust cross-modal systems for aerial object detection. To address this gap, our study proposes a brain-eye-computer based system for dim target detection in aerial images. This system detects the suspicious object region with the region proposal networks and employs eye-tracking technology to simultaneously extract the attention regions and EEG signal fragments of the subject from the images. These complementary signals are fused via an adaptive fusion framework, achieving enhanced detection precision, particularly for challenging dim targets.
@@ II-B Multimodal Learning
Modeling and analyzing data from various sensory modalities, including images, speech, text, and EEG signals, are key aspects of multimodal learning [25, 26]. Significant attention has been garnered by multimodal learning in the fields of artificial intelligence and machine learning in recent years, resulting in remarkable progress [9, 27].
Due to the development of deep learning, researchers propose novel loss functions or training approaches to optimize multimodal models. For instance, Du et al. [25] modeled the relationships between brain, visual, and linguistic features using multimodal deep generative models. They maximized both intra-modality and inter-modality mutual information regularization terms. Their approach addresses limitations such as the under-exploitation of multimodal knowledge and the scarcity of training data. Yang et al. [12] proposed a multimodal fusion approach for remote sensing image-audio retrieval tasks. They converted audio inputs to text and fused them with the text information to obtain a fusion representation. They optimized the common retrieval space using triplet loss, semantic loss, and consistency loss. Experimental findings on multiple datasets indicate the effectiveness of their approach.
However, these methods focus on feature fusion and loss function design when developing multimodal fusion algorithms, neglecting the potential benefits of knowledge transfer and mutual learning between different modalities. In this paper, we propose a dynamic hybrid fusion approach that integrates multimodal features and trains all branches in the model using an online knowledge distillation (OKD) method. This mutual learning approach for multimodal fusion leverages the visual and cognitive domains to effectively extract crucial information.
@@ II-C Knowledge Distillation
Knowledge Distillation (KD) has attracted considerable attention because of its capacity to compress models. This technique involves training a small student model by imitating the output distribution of a large teacher model [13]. KD has been extensively employed in different fields, including computer vision [28], natural language processing [29], and emotion recognition [30]. These approaches can be classified into offline distillation and online distillation according to the distillation schemes [31]. Offline distillation typically follows a two-stage training process where a trained teacher model guides the training of a student model. However, this approach is expensive and knowledge transfer is unidirectional. In contrast, online approaches enable end-to-end training, allowing for simultaneous learning of teacher and student models while reducing training [32].
@@
Current online knowledge distillation approaches have achieved significant progress with multimodal homogeneous data. For instance, Zhang et al. [32] introduced DML, demonstrating that student models can learn from each other through their predictions in deep mutual learning. Li et al. [35] proposed an innovative embedded OKD approach to leverage ensemble information, overall feature representations from peer networks, and logits to fully exploit the potential of networks. Gou et al. [36] proposed CKD-MKT, a collaborative knowledge distillation framework that integrates self-learning and online distillation for multitype knowledge transfer, enabling students to learn from both individual instances and instance relations with two peer networks.
However, offline distillation follows a one-way training paradigm, resulting in high computational costs. Most existing OKD methods are designed for homogeneous data and lack targeted fusion strategies for heterogeneous data. To address this limitation, we propose the AMBOKD method, which effectively extracts features from both 2D EEG data and 3D image data while treating the fusion model’s output features output as a single modality for mutual learning. Our method adaptively adjusts the influence weights and optimization levels of each modality, leading to more effective parameters optimization and superior fusion model performance.
@@ III System Design and Data Collection
In this section, we start by introducing the structure and data acquisition process of the brain-eye-computer based object detection system. Then, we introduce the designed ESSVP paradigm and its experimental environment in detail. Lastly, we present the preprocessing process and the details of the ESSVP dataset.